Hannah with a mass of 56.2 kg) is traveling at a speed of 12.8 m/s at the top of a 19.5-m high roller coaster Loop in a THEME park.


Find the following:

A.Hannah’s kinetic energy at the top of the loop.

B.Hannah’s potential energy at the top of the loop.

C.Assuming negligible losses of energy due to friction and air resistance, determine Hannah’s total

mechanical energy at the bottom of the loop (h=0 m).

D.Determine Hannah’s speed at the bottom of the loop.


1
Expert's answer
2022-03-01T15:32:54-0500

A.

"E_K=\frac{mv^2}2=4.6~kJ,"

B.

"E_P=mgh=10.7~kJ,"

C.

"E=E_P+E_K=15.3~kJ,"

D.

"v_1=v+\sqrt{2gh}=32.3~\frac ms ."